What percent discount would you have gotten if the actual price of an item was $16.95 and you paid $15.96?
Found 2 solutions by Maths68, richwmiller:
Answer by Maths68(1474)

Actual Amount = 16.95
Paid Amount = 15.96
Difference = 0.99
% discount = (Difference/Actual Amount)*100
% discount = (0.99/16.95)*100
% discount = (0.0584)*100
% discount = 5.84


Answer by richwmiller(17219)

1695-1596=99
99/1695*100=5.84%